OpenStudy (anonymous):

Please help! A water tank in the form of an inverted frustum of a cone has an altitude of 8 ft., and upper and lower radii of 6 ft. and 4 ft., respectively. Find the volume of the water tank and the wetted part of the tank if the depth of the water is 5 ft.

OpenStudy (aaronandyson):

Do you have any options given?

OpenStudy (aaronandyson):

When the depth is 5 ft, the tank is 5/8 full, so the top radius of the water is (4 ft) + (5/8)(6 ft - 4 ft) = 4 ft + (5/8)(2 ft) = 4 ft + 1.25 ft = 5.25 ft. The slant height 5.154 ft is the square root of [(5 ft)^2 + (5.25ft - 4ft)^2] = sqrt(25 ft^2 + 1.5625 ft^2) = 5.15388... feet A formula for the surface area of the frustum (see Wikipedia, etc) results in A = pi(4.00 ft + 5.25 ft)*5.154 ft + pi*(4.00 ft)^2 = 200.0 ft^2 source google and yahoo